OpenStudy (anonymous):

What number must you add to complete the square? x^2+8x=3

OpenStudy (anonymous):

If you divide 8 by 2, you get what?

OpenStudy (anonymous):

@skay 4

OpenStudy (anonymous):

yup. so then if you factor that you get (x-4)^2 which covers your x squared but whats this, 4^2 isnt 8....

OpenStudy (anonymous):

@skay 4^2 is 16

OpenStudy (anonymous):

very good. now, what can you do with that 16.

OpenStudy (anonymous):

@skay add 16

OpenStudy (anonymous):

lol yea. but remember you have to add 16 to both sides. what you do to one side you have to do to the other,.

OpenStudy (anonymous):

oh i made a mistake its (x+4)^2

OpenStudy (anonymous):

also! just to make sure, you should have x^2+8x+16=3+16 (x+4)^2=19 sqrt(x-4)^2)=sqrt(19) x-4=sqrt(19) x=4sqrt(19) that's if you're solving for x though x^2+8x+16 is a complete square.

OpenStudy (anonymous):

you can use this method for completing any square. hope it helped!